OpenStudy (anonymous):

Find the exact value. Use a sum or difference identity. sin(-105˙)

OpenStudy (mattfeury):

-105 degrees = 255 degrees 255˙ - 180˙= 75˙, so this is a 75 degree angle in the 3rd quadrant of the unit circle, meaning x and y are both negative.

OpenStudy (anonymous):

Omg, thank you so muchh! :D

OpenStudy (mattfeury):

no problem. please give medals to answerers that are helpful :) so sin(-105˙) = - sin(75˙)

OpenStudy (anonymous):

Thankss!!!

OpenStudy (anonymous):

Wait, but then what is sin(75˙) ?

OpenStudy (mattfeury):

you can use the half angle identity, just like here: http://www.schooltrainer.com/homework-answers/use-a-half-angle-identity-to-find-the-exact-value-sin75-degrees

OpenStudy (anonymous):

That helps but I have to use the Angle Sum/Difference Identities.

OpenStudy (anonymous):

Don't know how Mathematica did it but here is what you will have to shoot for: \[\text{Sin}[-105\text{Degree}]=-\frac{1+\sqrt{3}}{2 \sqrt{2}} \]